sin 3pi/8 cos pi/8 + cos3pi/8 sin pi/8
Answer by lwsshak3(11628)

sin 3pi/8 cos pi/8 + cos3pi/8 sin pi/8
Trig identity: sin(s+t)=sin s cos t + cos s sin t
For given problem:
sin 3π/8 cos π/8 + cos3 π/8 sin π/8
=sin(3π/8+π/8)
=sin(4π/8)
=sin(π/2)
=1